# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 2.3. Variational formulas of generalized angle
(ii): Suppose $c_{ijk}$ is space-like, then $||c_{ijk}||^{2}=1$ and $v_{j}*c_{ijk}=\pm\cosh q_{j}$, thus $||v_{j}\otimes c_{ijk}||=\sinh q_{j}$ if $q_{j}>0$ and $||v_{j}\otimes c_{ijk}||=-\sinh q_{j}$ if $q_{j}<0$ by (25).
If $c_{ijk}$ is in one of the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{V}},D_{\mathrm{VI}}$, then $q_{j}>0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$ by Lemma 2.13. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{III}}$, then $q_{j}<0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ by Lemma 2.13. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{IV}}$, then $q_{j}>0$ and $h_{k}<0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$ by Lemma 2.13. Note that $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},(c_{ijk})^{\bot}))=-\sinh(-h_{k})=\sinh h_{k}$.
If $c_{ijk}$ is in one of the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{vi}}$, then $q_{j}>0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$ by Lemma 2.15. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{iii}}$, then $q_{j}<0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ Lemma 2.15. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{v}}$, then $q_{j}>0$ and $h_{k}<0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$ by Lemma 2.15. And $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{iv}}$, then $q_{j}<0$ and $h_{k}<0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=-\cosh h_{k}$ by Lemma 2.15. And $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
In summary, if $c_{ijk}$ is space-like, then
@@
(iii): The case that $c_{ijk}$ is light-like follows from continuity by moving the center to the light cone. In fact, if $c_{ijk}$ is light-like, then $||c_{ijk}||^{2}=0$. Combining (28) and (26) gives
$\frac{\sinh^{2}\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||^{2}}{(c_{ij}\ast c_{ijk})^{2}}=\frac{||c_{ij}\otimes c_{ijk}||^{2}}{(c_{ij}\ast c_{ijk})^{2}}=1.$
Since $||c_{ij}\otimes c_{ijk}||=-(c_{ij}\ast c_{ijk})$, then
@@
Actually, the formulas in the proof of Lemma 2.18 are the sine and cosine laws of generalized right-angled hyperbolic triangles. In order to avoid discussing one by one, we use Lorentzian inner product and Lorentzian cross product to unify calculations.
@@
Under the same notations as in the proof of Lemma 2.18, a routine calculation gives
@@
The formula (29) was first obtained by Glickenstein-Thomas [3] for generalized discrete conformal structures on closed surfaces, which has lots of applications. See [9, 10, 12, 13, 14, 15] and others for example. In [11], Xu showed that the formula (29) holds for a special class of discrete conformal structures on surfaces with boundary and conjectured this formula holds for generalized discrete conformal structures on surfaces with boundary. Lemma 2.20 provides an affirmative answer to Xu’s conjecture.
Since the variational formula of generalized angle (15) is not symmetric in $i$ and $j$, we need to obtain a symmetric variational formula of generalized angle by reparameterization.
@@
There is a change of variables $u_{i}=u_{i}(f_{i})$ such that
@@
The function $u$ is also called a discrete conformal factor.
@@
Take new coordinates $u_{i}=u_{i}(f_{i})$ such that $\partial\theta^{jk}_{i}/\partial u_{j}=\partial\theta^{ik}_{j}/\partial u_{i}$. By (15), we have
@@
Recall the following formulas from the proof of [16, Theorem 1.6]
@@
By Remark 1.7, we can set $e^{C_{ij}}=1$. If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{f_{i}-f_{j}}$, then
@@
Suppose $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{f_{i}-f_{j}}$. If $d_{ij}<0$, then $d_{ik}<0$ by Lemma 2.17, and thus $d_{ji}>0,\ d_{ki}>0$. We need $d_{jk}>0,\ d_{kj}>0$ to make the condition (1) hold. Hence $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-e^{f_{i}-f_{k}}$ and $\frac{\sinh d_{jk}}{\sinh d_{ki}}=e^{f_{j}-f_{k}}$. Then
@@
If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, then
@@
By the arguments similar to that in the case $\mathrm{(i)}$, if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, then when $d_{ij}<0$,
$\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}};$
@@
$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=-\sqrt{1+\alpha_{j}e^{2f_{j}}}.$
@@
In fact, the function $u_{i}=u_{i}(f_{i})$ can be computed explicitly. If $\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}}$, then $u_{i}=-e^{-f_{i}}$. If $\frac{\partial f_{i}}{\partial u_{i}}=-e^{f_{i}}$, then $u_{i}=e^{-f_{i}}$.
@@
Specially, if $\alpha_{i}=-1$, then
@@
which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Note that $\sqrt{1-e^{2f_{i}}}\neq 0$ and thus $f_{i}\neq 0$. Then $f_{i},u_{i}\in\mathbb{R}_{<0}$. If $\alpha_{i}=1$, then
@@
which implies $e^{f_{i}}=-\frac{1}{\sinh u_{i}}$ and $f_{i}\in\mathbb{R}$, $u_{i}\in\mathbb{R}_{<0}$.
Suppose $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}$, then if $\alpha_{i}=0$, then $u_{i}=-f_{i}\in\mathbb{R}$; if $\alpha_{i}\neq 0$, then
@@
Specially, if $\alpha_{i}=-1$, then
@@
which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Similarly, the case that $f_{i}=0$ is removed. Then $f_{i}\in\mathbb{R}_{<0},\ u_{i}\in\mathbb{R}_{>0}$. If $\alpha_{i}=1$, then
@@
which implies $e^{f_{i}}=\frac{1}{\sinh u_{i}}$ and $f_{i}\in\mathbb{R},\ u_{i}\in\mathbb{R}_{>0}$.
@@ 3.1. Admissible space of the discrete conformal structure (A1)
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}^{E}_{>0}$. The admissible space of the discrete conformal factors $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (5) is defined to be
(33) $\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|$ $\cosh l_{rs}=-\sqrt{(1+\alpha_{r}e^{2f_{r}})(1+\alpha_{s}e^{2f_{s}})}$
@@
By a change of variables $u_{i}=u_{i}(f_{i})$ in Lemma 2.22, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transferred to the following admissible space $\mathcal{U}_{ijk}(\eta)$ of the discrete conformal factors $u$ for a right-angled hyperbolic hexagon $\{ijk\}$
$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}|(f_{i},f_{j},f_{k})\in\Omega_{ijk}(\eta)\},$
@@
Since $\frac{\sinh d_{rs}}{\sinh d_{sr}}=\sqrt{\frac{1+\alpha_{r}e^{2f_{r}}}{1+\alpha_{s}e^{2f_{s}}}}>0$ in Theorem 1.6, then by the proof of Lemma 2.22, we have
@@
According to Remark 2.23, we can describe the admissible space $\mathcal{U}_{ijk}(\eta)$ explicitly as follows:
@@
If $\alpha\equiv-1$, then $\cosh l_{rs}=-\sqrt{(1-e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\tanh u_{r}\tanh u_{s}+\eta_{rs}\frac{1}{\cosh u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $\cosh(u_{r}+u_{s})<\eta_{rs}$. Since $u\in\mathbb{R}^{3}_{<0}$, then $u_{r}+u_{s}>-\mathrm{arccosh}\ \eta_{rs}$. It is worth noting that $\eta_{rs}>1$.
@@
If $\alpha\equiv 1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1+e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\coth u_{s}+\eta_{rs}\frac{1}{\sinh u_{r}}\frac{1}{\sinh u_{s}}>1$, which implies $\cosh(u_{r}+u_{s})<\eta_{rs}$. Since $u\in\mathbb{R}^{3}_{<0}$, then $u_{r}+u_{s}>-\mathrm{arccosh}\ \eta_{rs}$. It is worth noting that $\eta_{rs}>1$.
@@
If $\alpha_{r}=0,\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{1-e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\tanh u_{s}+\eta_{rs}e^{u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
@@
If $\alpha_{r}=0,\alpha_{s}=1$, then $\cosh l_{rs}=-\sqrt{1+e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\coth u_{s}+\eta_{rs}e^{u_{r}}\frac{-1}{\sinh u_{s}}>1$. Since $u_{s}\in\mathbb{R}_{<0}$, then $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.

### v3

@@ 2.3. Variational formulas of generalized angle
(ii): Assume that $c_{ijk}$ is space-like. Then $\|c_{ijk}\|^{2}=1$ and $v_{j}*c_{ijk}=\pm\cosh q_{j}$. It follows from (24) that $\|v_{j}\otimes c_{ijk}\|=\sinh q_{j}$ if $q_{j}>0$, and $\|v_{j}\otimes c_{ijk}\|=-\sinh q_{j}$ if $q_{j}<0$. We proceed to analyze different cases based on the domain to which $c_{ijk}$ belongs.
If $c_{ijk}\in D_{\mathrm{I}}\cup D_{\mathrm{II}}\cup D_{\mathrm{V}}\cup D_{\mathrm{VI}}$, then Table 1 implies $q_{j}>0$ and $h_{k}>0$. By Lemma 2.12, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$. Additionally, it is known that $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}\in D_{\mathrm{III}}$, then Table 1 implies $q_{j}<0$ and $h_{k}>0$. By Lemma 2.12, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ by Lemma 2.12. Additionally, it is known that $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}\in D_{\mathrm{IV}}$, then Table 1 implies $q_{j}>0$ and $h_{k}<0$. By Lemma 2.12, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$. Note that $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},(c_{ijk})^{\bot}))=-\sinh(-h_{k})=\sinh h_{k}$.
If $c_{ijk}\in D_{\mathrm{i}}\cup D_{\mathrm{ii}}\cup D_{\mathrm{vi}}$, then Table 1 implies $q_{j}>0$ and $h_{k}>0$. By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$. Additionally, $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}\in D_{\mathrm{iii}}$, then Table 1 implies $q_{j}<0$ and $h_{k}>0$. By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$. Additionally, $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}\in D_{\mathrm{v}}$, then Table 1 implies $q_{j}>0$ and $h_{k}<0$. By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$. Additionally, $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
If $c_{ijk}\in D_{\mathrm{iv}}$, then Table 1 implies $q_{j}<0$ and $h_{k}<0$. By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot(-\sinh q_{j})=-\cosh h_{k}$. Additionally, $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
In all the above cases for space-like $c_{ijk}$, we can conclude that
@@
(iii): The case where $c_{ijk}$ is light-like can be derived by continuity, specifically by considering the limit as the face center approaches the light cone. In detail, for a light-like $c_{ijk}$, we have $\|c_{ijk}\|^{2}=0$. Combining (27) and (25) gives
$\frac{\sinh^{2}\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|^{2}}{(c_{ij}\ast c_{ijk})^{2}}=\frac{\|c_{ij}\otimes c_{ijk}\|^{2}}{(c_{ij}\ast c_{ijk})^{2}}=1.$
Furthermore, since $\|c_{ij}\otimes c_{ijk}\|=-(c_{ij}\ast c_{ijk})$, it follows that
@@
The formulas derived in the proof of Lemma 2.17 essentially correspond to the sine and cosine laws for generalized right-angled hyperbolic triangles. To avoid the need for cumbersome case-by-case analyses, we adopt the Lorentzian inner product and Lorentzian cross product as a unifying framework for the calculations.
@@
Under the same notations as used in the proof of Lemma 2.17, analogous calculations yield
@@
Formula (28) was first established by Glickenstein-Thomas [3] in the context of discrete conformal structures on closed surfaces, and it has since been widely applied. For instance, relevant discussions can be found in [9, 12, 11, 13] and related works. In [10], Xu showed that formula (28) remains valid for a specific class of discrete conformal structures on surfaces with boundary, and conjectured that this formula holds for generic discrete conformal structures on surfaces with boundary. Lemma 2.19 provides an affirmative resolution to Xu’s conjecture.
Given that the variational formula for the generalized angle (14) lacks symmetry in $i$ and $j$, we aim to derive a symmetric variational formula for the generalized angle via reparameterization.
@@
There exists a change of variables $u_{i}=u_{i}(f_{i})$ such that
@@
The function $u$ is also referred to as a discrete conformal factor.
@@
Suppose there exists a change of variables $u_{i}=u_{i}(f_{i})$ such that $\partial\theta^{jk}_{i}/\partial u_{j}=\partial\theta^{ik}_{j}/\partial u_{i}$. It follows from (14) that
@@
From the proof of [14, Theorem 1.5], we recall the identities
@@
By Remark 1.6, we set $e^{C_{ij}}=1$. If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{f_{i}-f_{j}}$, then
@@
Suppose $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{f_{i}-f_{j}}$. If $d_{ij}<0$, then $d_{ik}<0$ by Lemma 2.16, which implies $d_{ji}>0$ and $d_{ki}>0$. For the condition (2) to hold, we require $d_{jk}>0$ and $d_{kj}>0$. It follows that $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-e^{f_{i}-f_{k}}$ and $\frac{\sinh d_{jk}}{\sinh d_{ki}}=e^{f_{j}-f_{k}}$. Hence
@@
If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then
@@
If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then $\alpha\equiv-1$ and
$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{e^{2f_{j}}-1}.$
(iii):
If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then by arguments analogous to those in the case (i), when $d_{ij}<0$,
$\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}},\quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{1+\alpha_{k}e^{2f_{k}}};$
@@
$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}},\quad\frac{\partial f_{j}}{\partial u_{j}}=-\sqrt{1+\alpha_{j}e^{2f_{j}}}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{1+\alpha_{k}e^{2f_{k}}}.$
If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then when $d_{ij}<0$,
(30) $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{e^{2f_{i}}-1},\quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{e^{2f_{j}}-1}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{e^{2f_{k}}-1};$
when $d_{ji}<0$,
$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1},\quad\frac{\partial f_{j}}{\partial u_{j}}=-\sqrt{e^{2f_{j}}-1}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{e^{2f_{k}}-1}.$
@@
In fact, the function $u_{i}=u_{i}(f_{i})$ admits an explicit expression.
(i):
If $\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}}$, then $u_{i}=-e^{-f_{i}}$. If $\frac{\partial f_{i}}{\partial u_{i}}=-e^{f_{i}}$, then $u_{i}=e^{-f_{i}}$.
(ii):
@@
In particular, if $\alpha_{i}=-1$, then
@@
which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Note that $\sqrt{1-e^{2f_{i}}}\neq 0$ and thus $f_{i}\neq 0$. Consequently, $f_{i}\in\mathbb{R}_{<0}$ and $u_{i}\in\mathbb{R}_{<0}$. If $\alpha_{i}=1$, then
@@
which implies $e^{f_{i}}=-\frac{1}{\sinh u_{i}}$ with $f_{i}\in\mathbb{R}$ and $u_{i}\in\mathbb{R}_{<0}$.
(iii):
Suppose $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}$, if $\alpha_{i}=0$, then $u_{i}=-f_{i}\in\mathbb{R}$; if $\alpha_{i}\neq 0$, then
@@
In particular, if $\alpha_{i}=-1$, then
@@
which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Similarly, the case that $f_{i}=0$ is excluded, so $f_{i}\in\mathbb{R}_{<0}$ and $u_{i}\in\mathbb{R}_{>0}$. If $\alpha_{i}=1$, then
@@
which implies $e^{f_{i}}=\frac{1}{\sinh u_{i}}$ with $f_{i}\in\mathbb{R}$ and $u_{i}\in\mathbb{R}_{>0}$.
(iv):
If $\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1}$, then
$\tan u_{i}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(0,\frac{\pi}{2}),$
$-\frac{1}{\tan u_{i}}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(-\frac{\pi}{2},0).$
This implies that $e^{f_{i}}=\frac{1}{\cos u_{i}}$ for $u_{i}\in(0,\frac{\pi}{2})$, and $e^{f_{i}}=-\frac{1}{\sin u_{i}}$ for $u_{i}\in(-\frac{\pi}{2},0)$. If $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{e^{2f_{i}}-1}$, then
$-\tan u_{i}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(-\frac{\pi}{2},0),$
$\frac{1}{\tan u_{i}}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(0,\frac{\pi}{2}).$
This implies that $e^{f_{i}}=\frac{1}{\cos u_{i}}$ for $u_{i}\in(-\frac{\pi}{2},0)$, and $e^{f_{i}}=\frac{1}{\sin u_{i}}$ for $u_{i}\in(0,\frac{\pi}{2})$.
@@ 3.1. Admissible space of the discrete conformal structure (A1)
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weights are given by $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}^{E}_{>0}$. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (5), the admissible space of the discrete conformal factors $f$ is defined as
(33) $\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid$ $\cosh l_{rs}=-\sqrt{(1+\alpha_{r}e^{2f_{r}})(1+\alpha_{s}e^{2f_{s}})}$
@@
By a change of variables $u_{i}=u_{i}(f_{i})$ in Lemma 2.21, the admissible space $\Omega_{ijk}(\eta)$ is transformed into the following admissible space $\mathcal{U}_{ijk}(\eta)$ of the discrete conformal factors $u$ for a right-angled hyperbolic hexagon $\{ijk\}$:
$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}\mid(f_{i},f_{j},f_{k})\in\Omega_{ijk}(\eta)\},$
@@
Since $\frac{\sinh d_{rs}}{\sinh d_{sr}}=\sqrt{\frac{1+\alpha_{r}e^{2f_{r}}}{1+\alpha_{s}e^{2f_{s}}}}>0$ with $01+\alpha_{r}e^{2f_{r}}>$ 0 and $01+\alpha_{s}e^{2f_{s}}>$ 0 in Theorem 1.5, it follows from the proof of Lemma 2.21 that
@@
By Remark 2.22, the following results are obtained:
@@
If $\alpha\equiv-1$, then $\cosh l_{rs}=-\sqrt{(1-e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\tanh u_{r}\tanh u_{s}+\eta_{rs}\frac{1}{\cosh u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $\cosh\,(u_{r}+u_{s})<\eta_{rs}$. Since $u\in\mathbb{R}^{3}_{<0}$, it follows that $u_{r}+u_{s}>-\mathrm{arccosh}\,\eta_{rs}$. Note that $\eta_{rs}>1$.
@@
If $\alpha\equiv 1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1+e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\coth u_{s}+\eta_{rs}\frac{1}{\sinh u_{r}}\frac{1}{\sinh u_{s}}>1$, which implies $\cosh\,(u_{r}+u_{s})<\eta_{rs}$. Since $u\in\mathbb{R}^{3}_{<0}$, it follows that $u_{r}+u_{s}>-\mathrm{arccosh}\,\eta_{rs}$. Note that $\eta_{rs}>1$.
@@
If $\alpha_{r}=0$ and $\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{1-e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\tanh u_{s}+\eta_{rs}e^{u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
@@
If $\alpha_{r}=0$ and $\alpha_{s}=1$, then $\cosh l_{rs}=-\sqrt{1+e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\coth u_{s}+\eta_{rs}e^{u_{r}}\frac{-1}{\sinh u_{s}}>1$. Since $u_{s}\in\mathbb{R}_{<0}$, it follows that $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.